OpenStudy (anonymous):

Solve sin2x - 2cos^2x=0

OpenStudy (haseeb96):

2sinxcosx -2 (1-sin^2 x)=0 2(sinxcosx-1+sin^2x)=0 2(sinxcosx+sin^2x)-2=0 2sinx(cosx+sinx)=+2 sinx(cosx+sinx)=1 cosx+sinx=1\sinx cosx=1\sinx - sinx cosx=1-sin^2x\sinx cosx=cos^2 x \sinx sinx=cosx